# Fast-weight Product Key Memory

### v1

@@ 4.2 PPL Evaluation
We evaluate perplexity (PPL) on three distinct datasets to assess different memory capabilities: Fineweb-Edu (knowledge-intensive, short context), LC64 (in-domain, long context), and LAMBADA (LAMBADA, out-of-domain, long context). Each evaluation dataset contains 8M tokens. We evaluate 4K-token sequences in their original order with a batch size of 1 so that sequences from the same document are presented sequentially. This enables fast weights to capture the dependency between adjacent sequences. Figure 2 shows the perplexities on the datasets for all models.
Finding 1: FwPKM and PKM serve distinct, complementary roles
Baseline models with layouts GDN and GDN+SWA lack components for modeling long-range dependency, and their FwPKM versions address the weakness. As shown in Figure 2, FwPKM significantly reduces PPL on long-context datasets (LC64, LAMBADA), confirming its role as an episodic memory. In contrast, standard PKM provides the largest gains on Fineweb-Edu, acting as a semantic memory for global knowledge. The combination of both modules yields the best performance across all metrics, suggesting they address orthogonal limitations in the baseline architectures.
Figure 2: Perplexity on Fineweb-Edu, LC64, and LAMBADA
Figure 3: FwPKM gating value distribution on Fineweb-Edu, LC64, and LAMBADA test sets. Each row represents one FwPKM layer.
Finding 2: FwPKM competes with Full Attention
@@
To mitigate the above issue, we restrict full attention’s long-range perception at training time, by imposing a sliding attention window of length 512 to all full attentions with a probability of 0.9. Denoted by a suffix pSWA, these new models learn to use FwPKM more as suggested by more gating value distribution mass on the high-value end in Figure 3. While pSWA has minimal impact in the PPL evaluation, we will see a more significant difference in the following section.
4.3 NIAH Evaluation
We conduct Needle in a Haystack (NIAH, NIAH; LandmarkAttention) evaluation to further verify FwPKM’s functionality as episodic memory.
@@ 4.3 NIAH Evaluation
Iterative Memorization ($n$-iter NIAH) A unique feature of FwPKM is its ability to improve memory fidelity by re-processing the same input. We define a $n$-iter NIAH setting, where the model forwards the same haystack context $n$ times before answering. FwPKM’s process chunk size $C$ is accordingly changed to the length of each haystack context such that it updates its memory once after reading an entire haystack. The basic setting is $1$-iter NIAH.
@@
For GDN and GDN+SWA layouts, a single pass ($1$-iter) is often insufficient for perfect retrieval. However, a second pass ($2$-iter) yields a massive boost in accuracy (jumping from $10\%$ to $70\%$ in many cases). This confirms that FwPKM effectively exploits test-time training to consolidate episodic memories, surpassing softmax attention. In addition, effective iterative memorization (in Figure 4) correlates with high gating values (in Figure 3).
@@
Finding 6: Longer context requires more memorization iterations.
As context length grows from 4K to 128K, the $2$-iter accuracy significantly reduces from $70\%$ to $40\%$ for GDN | PKM@6 + FwPKM@2,10 and more severely for other models. However, more iterations largely fill in the gap.
@@ 5.1 Case Study: Probing Memory in NIAH
Figure 5: An example of FwPKM slot access of GDN+PKM@6+FwPKM@2,10 during generating an NIAH-4K answer. The model memorizes the haystack for 3 extra iterations, i.e. 44-iter NIAH.
We analyze the memory access patterns of the GDN+PKM@6+FwPKM@2,10 model during the $4$-iter NIAH task (4K context). At each generation step, every memory layer returns Top-$8$ retrieved slots. We show the content stored in each slot from the latest memorization [^3] A slot may contain information from earlier updates, but we omit them due to space limit., namely the query token, the target value token, and its surrounding context. In addition, a [HIT] label indicates whether a slot stores the ground-truth next token.
@@
As shown in Figure 5, the majority of retrieved slots contain the correct target tokens associated with the query needle.
Error analysis
Incorrect retrievals typically share the same query token as the input but lack the correct prefix context. This suggests that retrieval errors stem from imperfect prefix matching in the latent query space rather than storage failure.
@@
Despite the slot errors, the model successfully aggregates information across the $16$ slots from 2 FwPKM layers to generate the correct 6-digit value, demonstrating robust distributed storage.
@@ 5.2 Case Study: Selective Gating
We further examine the “gating values” $g_{t}$ to understand when the model chooses to rely on episodic memory versus its static slow weights. Using the Wikipedia article for “Sakana AI” [^4] https://en.wikipedia.org/wiki/Sakana_AI as a test case, we change the model’s update chunk size from 512 to 32 to adapt for the article’s length. As shown in Figure 6, we can observe distinct behaviors across layers.
@@
The lower-layer FwPKM tends to maintain high gating values across all tokens, serving as a general-purpose buffer. In contrast, the higher-layer FwPKM exhibits highly selective activation.
@@
In the higher layer, gating values spike specifically for tokens related to rare named entities (e.g.Sakana AI”, “David Ha”, “Llion Jones”, and “Ren Ito”). This indicates that the model effectively distinguishes between general linguistic patterns and novel, context-specific entities, which are processed by slow weights and fast weights, respectively.
Figure 6: GDN+PKM@6+FwPKM@2,10’s FwPKM gating values on tokens from the Wikipedia article for Sakana AI.
@@ 6 Cost Analyses
We compare the main models’ parameter numbers and computation costs in Table 2. In particular, we report FLOPs (Floating Point Operations) to measure the required computation as well as FLOPS (Floating Point Operations Per Second), which is FLOPs divided by running time (in seconds), to additionally take into account implementation efficiency.
The sparsity of PKM/FwPKM makes it even less FLOPs-intensive than a baseline MLP layer despite much increased model size. However FLOPS numbers suggest that PKM/FwPKM components take longer to run. One main reason is the large gap between their efficient implementation. Softmax attention and linear attention (e.g. GDN) are fast using kernels from FlashAttention (FlashAttention; FlashAttention2) and FlashLinearAttention (FLA). An important future direction for this work is to design more efficient kernels to facilitate easier scaling up and a broader adoption of the FwPKM model.
Table 2: Comparison of model size and computation cost.
@@
GDN 112.11 22.05 121.82 42.81
     $|$ PKM@2,6,10 509.08 20.91 84.47 30.09
     $|$ FwPKM@2,6,10 519.12 22.74 46.44 16.10
     $|$ PKM@6 + FwPKM@2,10 515.77 22.14 54.63 19.26
GDN+SWA 112.65 22.14 124.42 43.45
     $|$ PKM@2,6,10 509.61 21.03 88.16 31.25
     $|$ FwPKM@2,6,10 519.66 22.86 47.16 16.27
     $|$ PKM@6 + FwPKM@2,10 516.31 22.23 55.30 19.39
GDN+FA 112.65 22.14 120.67 42.08
     $|$ PKM@2,6,10 509.61 21.03 84.69 30.07
     $|$ FwPKM@2,6,10 519.66 22.86 46.51 16.05
     $|$ PKM@6 + FwPKM@2,10 516.31 22.23 54.99 19.29
FA 114.25 22.47 120.45 41.55
     $|$ PKM@2,6,10 511.22 21.33 81.22 28.49
     $|$ FwPKM@2,6,10 521.26 23.16 46.63 15.89
     $|$ PKM@6 + FwPKM@2,10 517.91 22.56 54.43 18.84
@@ 7 Related Work
Hybrid architectures Recognizing the complementary strengths of different sequence models, recent work has increasingly focused on hybrid architectures that combine the high-fidelity retrieval of quadratic attention with the efficiency of linear or recurrent layers. This includes hybrid models trained from scratch (HybridFAandDeltaNet), as well as large-scale Hybrid LLMs such as Samba (Samba) and KimiLinear (KimiLinear). QwenNext also employs this interleaved design. Furthermore, approaches like Artificial Hippocampus Networks (AHN, AHN) explore fine-tuning techniques to integrate these distinct memory systems effectively. Experiments in Section 4 demonstrated interactions between memories of different characteristics. The combination of linear attention (i.e. GDN), softmax attention, slow-weight sparse memory (i.e. PKM), and fast-weight sparse memory (i.e. FwPKM) implements a versatile memory system that excels at various tasks.
@@ 8 Conclusion
In this work, we introduced Fast-weight Product Key Memory (FwPKM), a memory-augmented layer that unifies PKM’s large-scale sparse storage with the rapid adaptability of fast weights. Concretely, FwPKM extends Product Key Memory from a static retrieval module into a context-responsive component whose parameters can be updated online, allowing the model to write information into memory and later retrieve it. This addresses a key limitation of prior sparse memory modules, namely their inability to efficiently incorporate new evidence at inference time, and makes PKM better suited for long-context settings where relevant information may be separated by many thousands of tokens.
Empirically, we find that FwPKM remains effective far beyond its training regime. Models trained on 4K-token sequences generalize to 128K-token contexts, while exhibiting behavior consistent with a robust episodic memory that complements the semantic knowledge stored in standard layers.
At the same time, several challenges remain. Online updates introduce additional computation and choices such as chunk size, update frequency, and optimization hyperparameters. Scaling these updates efficiently motivates further systems work, including faster sparse update kernels and better implementation strategies. On the modeling side, future works include more powerful and robust memory design regarding its architecture, update rule, and retention. Overall, by integrating sparse storage with fast, context-driven updates, FwPKM offers a promising step toward language models with versatile and mutually complementary memory components.
@@ Acknowledgments
We thank Kai Arulkumaran, Luke Darlow, Stefania Druga and the rest of the Sakana AI team for fruitful discussions throughout the project.
Appendix A Detailed Training Settings
FwMLP
To rule out the impact of implementation details unrelated to the PKM architecture, we propose a baseline model that replaces the PKM in FwPKM with a SwiGLU-MLP that maintains three fast-weight matrices (and their biases) for up, gating, and down projection. The baseline, denoted as FwMLP, updates its fast weights by minimizing the MSE loss between its predicted values and target lookahead values at a chunk level. Due to its dense nature, we reduce the MSE losses in a chunk by averaging over both sample and feature dimensions, and we do not apply the loss aggregation and gradient shaping techniques mentioned in Section 3.2.2. For the same reason, addressing optimization in Section 3.3 is irrelevant too.
LaCT
We adopt the official implementation of LaCT (TTTPlus) [^5] https://github.com/a1600012888/LaCT/tree/main/lact_llm as a strong TTT (TTT) baseline. The LaCT architecture consists of a sliding window attention, a fast-weight SwiGLU MLP, and a slow-weight SwiGLU MLP in every layer. The fast weights are optimized to minimize a dot product loss via SGD with momentum or Muon (muon), and we found SGD with momentum achieves better performance in our experiments. LaCT uses data-dependent learning rate and L2 weight normalization to improve memorization and retention. One notable difference between LaCT (or more generally TTT) and FwMLP/FwPKM is that LaCT/TTT maintains an individual set of fast weights for each sequence in a mini batch, while FwMLP/FwPKM uses a shared set of fast weights for all sequences. We compared several configurations of sliding window size (W) and update chunk size (C). Among $512C+512W$, $512C+2048W$, $512C+4096W$, and $2048C+2048W$, the best model is $512C+2048W$.

### v2

@@ 4.2 PPL Evaluation
Figure 2: Perplexity on Fineweb-Edu, LC64, and LAMBADA. We use bar colors to help distinguish between models with different types of memory components.
Figure 3: FwPKM gating value distribution on Fineweb-Edu, LC64, and LAMBADA test sets. Each row represents one FwPKM layer. The probability densities are categorized into 10 bins. A darker color suggests a higher probability density of a bin.
We evaluate perplexity (PPL) on three distinct datasets to assess different memory capabilities: Fineweb-Edu (knowledge-intensive, short context), LC64 (in-domain, long context), and LAMBADA (LAMBADA, out-of-domain, long context). Each evaluation dataset contains 8M tokens.
Evaluation protocol
For PPL evaluation, we evaluate 4K-token sequences in their original order with a batch size of 1 so that segments are presented sequentially without resetting fast weights between adjacent segments. This enables fast weights to capture dependencies across segment boundaries. Figure 2 shows perplexities for all models.
Finding 1: FwPKM and PKM serve distinct, complementary roles.
Baseline models with layouts GDN and GDN+SWA lack components for modeling long-range dependency, and their FwPKM versions address the weakness. As shown in Figure 2, FwPKM significantly reduces PPL on long-context datasets (LC64, LAMBADA), confirming its role as an episodic memory. In contrast, standard PKM provides the largest gains on Fineweb-Edu, acting as a semantic memory for general knowledge. The combination of both modules yields the best performance across all metrics, suggesting they address orthogonal limitations in the baseline architectures. In Appendix D.2 and E.1, we further demonstrate that FwPKM is likely to exploit cross-sequence dependency to reduce PPL on long-context datasets.
Finding 2: FwPKM competes with Full Attention.
@@
To mitigate the above issue, we restrict full attention’s long-range perception at training time by imposing a sliding attention window of length 512 on all full attention layers with probability 0.9. Denoted by a suffix pSWA, these models learn to use FwPKM more, as suggested by increased high-end mass in the gating distributions in Figure 3. While pSWA has minimal impact on PPL, we will see a more significant difference in the following experiments.
4.3 Needle-in-a-Haystack Evaluation
We conduct Needle-in-a-Haystack (NIAH, NIAH; LandmarkAttention) evaluation to further verify FwPKM’s functionality as episodic memory. In all NIAH experiments, we update fast weights after reading the full haystack (and after each rereading iteration, see below); we do not perform additional updates while generating the answer.
@@ 4.3 Needle-in-a-Haystack Evaluation
Iterative Memorization ($n$-iter NIAH). A unique feature of FwPKM is its ability to improve memory fidelity by re-processing the same input. We define a $n$-iter NIAH setting, where the model forwards the same haystack context $n$ times before answering. FwPKM’s process chunk size $C$ is accordingly changed to the length of each haystack context such that it updates its memory once after reading an entire haystack. The basic setting is $1$-iter NIAH.
@@
For GDN and GDN+SWA layouts, a single pass ($1$-iter) is often insufficient for perfect retrieval. However, a second pass ($2$-iter) yields a massive boost in accuracy (jumping from $10\%$ to $50\%$ in many cases). More passes further improve the accuracy to $70\%$. This confirms that FwPKM effectively exploits test-time training to consolidate episodic memories, surpassing softmax attention. In addition, effective iterative memorization (in Figure 4) correlates with high gating values (in Figure 3).
@@
Figure 5: Perplexity on five Longbench tasks. We use bar colors to help distinguish between models with different types of memory components.
4.4 Longbench Evaluation
In addition to the synthetic NIAH task, we assess long-context modeling on realistic data from Longbench (Longbench). Longbench contains 21 tasks whose average length ranges from 5K to 15K whitespace-separated words. The tasks are mostly question answering and summarization, but their generative nature makes it difficult for pretrained-only (non instruction-tuned) models to produce well-calibrated free-form answers. Instead of computing official metrics such as F1 score, we use perplexity for evaluation.
We select 2WikiMultihopQA, HotpotQA, MultiFieldQA-en, MuSiQue, and NarrativeQA, resulting in 950 examples with an average length of 17421 tokens and a maximum length of 81616 tokens. We only compute perplexity on answer tokens. For examples with multiple answer references, we use the first reference for perplexity computation. We update fast weights once after reading the entire context (i.e., no updates during answer generation), mirroring the $1$-iter NIAH protocol. We omit the data for LaCT because it performs poorly on the Longbench data (PPL >100>100).
Finding 6: FwPKM reduces answer-token perplexity on realistic long-context tasks.
As shown in Figure 5, FwPKM effectively reduces perplexity on these long-context tasks for GDN, GDN+SWA, and GDN+FA models where FwPKM is actively utilized.
4.5 Continual Learning Evaluation
[…]
Beyond long-context modeling, we study whether the large-scale fast weights in FwPKM adapt to new data domains. We devise an adapt-and-test experiment. From the Pile corpus (Pile), we choose six subsets DM Mathematics, FreeLaw, PhilPapers, PubMed Central, Ubuntu IRC, and USPTO Backgrounds to represent different professional domains. For each domain, we sample a 4M-token adapt set and a 4M-token eval set without overlap.
We run 6 adaptation stages sequentially (one per domain’s adapt set), during which fast weights are updated while slow weights are frozen. After each adaptation stage, we compute the model’s average negative log-likelihood (NLL) on all 6 eval sets and report changes relative to the initial NLL before the first stage. In Figure 6, domains correspond to colors, used both for the background shading of each adaptation stage and for the lines showing $\Delta\text{NLL}$.
Finding 7: FwPKM adapts quickly but struggles with retention across stages.
As expected, after each adaptation stage, the NLL of the corresponding domain decreases as the fast weights store related domain knowledge. However, these gains often do not persist across subsequent stages, suggesting that previously stored fast-weight knowledge is flushed and replaced by newer domains. This motivates developing a memory retention mechanism to realize long-term continual learning in future work.
@@ 5.1 Case Study: Probing Memory in NIAH
Figure 7: An example of FwPKM slot access of GDN+SWA|PKM@6+FwPKM@2,10 during generating an NIAH-4K answer. The model memorizes the haystack for 3 extra iterations, i.e. 44-iter NIAH. See Section 5.1 for the visual elements’ definitions.
We investigate the internal retrieval mechanism by visualizing the memory access patterns of the GDN+SWA|PKM@6+FwPKM@2,10 model during a 4-iter Needle-in-a-Haystack (NIAH) task with a 4K context. Specifically, we trace the model’s internal state as it generates the 6-digit passkey answer.
Figure 7 displays the Top-$8$ retrieved memory slots for both a lower layer (Layer 2) and a higher layer (Layer 10) at each generation step. For every retrieved slot, we decode and display the content stored during the latest memorization step: the query token (the key used for addressing), the target value token (the predicted next token), and the surrounding context (the text window surrounding the target in the original haystack). To facilitate analysis, we mark slots with a [HIT] label if they explicitly store the ground-truth next token required for the current generation step.
@@
As shown in Figure 7, the majority of retrieved slots contain the correct target tokens associated with the query needle. The model successfully surfaces the specific needle (ID “8320” at depth $89.66\%$) from thousands of tokens, populating the top slots with the correct 6-digit value sequence (“964876”).
@@
Despite individual slot errors, the model successfully aggregates information across the $16$ slots from the two FwPKM layers to generate the correct 6-digit value. This demonstrates that FwPKM functions as a robust distributed storage system, where consensus across multiple slots compensates for noise in individual retrievals.
@@ 5.2 Case Study: Selective Gating
Figure 8: GDN+SWA|PKM@6+FwPKM@2,10’s FwPKM gating values on tokens from the Wikipedia article for Sakana AI. The blue color in the background denotes the gating intensity of the FwPKM at layer 2 and the red color denotes the FwPKM at layer 10. A darker color represents higher intensity. Black vertical lines show positions where FwPKM’s fast weights are updated, which is every 32 tokens as we specified for this visualization example.
We further examine the “gating values” $g_{t}$ to understand when the model chooses to rely on episodic memory versus its static slow weights. To visualize this, we feed the Wikipedia article for “Sakana AI” [^2] https://en.wikipedia.org/wiki/Sakana_AI—a topic unseen during the model’s pre-training—into the network. We adjust the update chunk size from 512 to 32 to capture fine-grained local adaptivity.
Figure 8 presents a heatmap of the gating scalar $g_{t}$ with input text in the foreground. Higher intensity (represented by darker bars) indicates a value of $g_{t}$ closer to $1$, signifying a strong reliance on the fast-weight episodic memory. Conversely, lower intensity indicates reliance on the pre-trained slow weights.
@@
We observe distinct behaviors across model depths. The lower-layer FwPKM tends to maintain high gating values across almost all tokens, acting as a general-purpose buffer that indiscriminately caches recent history. In contrast, the higher-layer FwPKM exhibits highly selective activation.
@@
As seen in Figure 8 (layer 10), gating values spike specifically for tokens related to rare named entities and proper nouns (e.g., “Sakana AI”, “David Ha”, “Llion Jones”, and “Ren Ito”). This indicates that the model effectively distinguishes between general linguistic patterns (processed by slow weights) and novel, context-specific entities (processed by fast weights).
@@ 6 Cost Analyses
We evaluate the scalability and efficiency of FwPKM by profiling model size and computational costs, summarized in Table 1. Our analysis focuses on three key metrics:
  • Active Parameters: The subset of parameters actually used during a forward pass, highlighting the architectural sparsity.
  • FLOPs (Floating Point Operations): The theoretical computational volume required for each run.
  • FLOPS (FLOPs per Second): The empirical hardware utilization efficiency, derived from actual running time.
Profiling was conducted by running 100 training steps on a single H100 GPU with a global batch size of 8, using the configuration detailed in Section 4.1 and Appendix C.
Sparsity and Computational Volume
Despite a substantial increase in total parameter count (approx. $5\times$ larger than baselines), the Active Parameter count of FwPKM models remains comparable to the dense baselines (e.g., GDN: 112M vs. FwPKM: 116M). Consequently, the theoretical computational cost (FLOPs) is remarkably efficient; in some cases, the sparsity of PKM/FwPKM makes them even less FLOPs-intensive than standard dense MLP layers.
Hardware Utilization and Kernels
While theoretical costs are low, the wall-clock speed (Samples per sec.) and hardware utilization (FLOPS) reveal an implementation gap. FwPKM components currently run slower than their dense counterparts. This is primarily due to the maturity difference in kernel optimization: softmax attention and GDN benefit from highly optimized kernels like FlashAttention (FlashAttention; FlashAttention2) and FlashLinearAttention (FLA). In contrast, our sparse operations rely on less specialized implementations. Developing dedicated, hardware-aware kernels for sparse fast-weight updates is a critical future direction to bridge this gap and facilitate broader adoption.
Table 1: Comparison of model size and computation cost. Active parameter counts reflect the sparse layer specifications: PKM utilizes 128 active rows, while FwPKM utilizes 8 active rows per token.
@@
Total Active
GDN 112.11 112.11 22.04 150.00 54.44
| PKM@2,6,10 509.08 106.62 20.92 112.19 42.90
| FwPKM@2,6,10 519.12 116.48 22.74 54.12 19.04
| PKM@6 + FwPKM@2,10 515.77 113.19 22.14 58.85 21.27
GDN+SWA 112.65 112.65 22.15 151.11 54.58
| PKM@2,6,10 509.61 107.15 21.03 109.43 41.63
| FwPKM@2,6,10 519.66 117.02 22.85 51.70 18.10
| PKM@6 + FwPKM@2,10 516.31 113.73 22.24 58.55 21.06
GDN+FA 112.65 112.65 22.16 147.71 53.33
| PKM@2,6,10 509.61 107.15 21.04 102.76 39.07
| FwPKM@2,6,10 519.66 117.02 22.86 51.91 18.16
| PKM@6 + FwPKM@2,10 516.31 113.73 22.25 57.69 20.74
FA 114.25 114.25 22.51 158.75 56.42
| PKM@2,6,10 511.22 108.76 21.39 106.40 39.79
| FwPKM@2,6,10 521.26 118.62 23.21 54.34 18.73
| PKM@6 + FwPKM@2,10 517.91 115.33 22.60 61.80 21.87
GDN | FwMLP@2,6,10 118.05 118.05 23.21 134.04 46.21
GDN+SWA | FwMLP@2,6,10 118.58 118.58 23.31 128.40 44.06
GDN+FA | FwMLP@2,6,10 118.58 118.58 23.32 97.18 33.33
FA | FwMLP@2,6,10 120.19 120.19 23.67 118.85 40.16
LaCT 112.70 112.70 24.44 92.38 30.24
@@ 7 Related Work
Hybrid architectures Recognizing the complementary strengths of different sequence models, recent work has increasingly focused on hybrid architectures that combine the high-fidelity retrieval of quadratic attention with the efficiency of linear or recurrent layers. This includes hybrid models trained from scratch (HybridFAandDeltaNet), as well as large-scale Hybrid LLMs such as Samba (Samba) and KimiLinear (KimiLinear). QwenNext (qwen3technicalreport; Qwen3Next80BThinking) also employs this interleaved design. Furthermore, approaches like Artificial Hippocampus Networks (AHN, AHN) explore fine-tuning techniques to integrate these distinct memory systems effectively. Experiments in Section 4 demonstrated interactions between memories of different characteristics. The combination of linear attention (i.e. GDN), softmax attention, slow-weight sparse memory (i.e. PKM), and fast-weight sparse memory (i.e. FwPKM) implements a versatile memory system that excels at various tasks.
@@ 8 Conclusion
In this work, we introduced Fast-weight Product Key Memory (FwPKM), a novel architecture that resolves the tension between memory capacity and computational efficiency in sequence modeling. By unifying the large-scale sparse storage of Product Key Memory (PKM) with the rapid adaptability of Test-Time Training (TTT), we transform PKM from a static retrieval module into a dynamic, context-responsive component. FwPKM performs Sparse TTT – updating key-value parameters online via sparse gradient descent – and enables models to efficiently memorize vast amounts of transient information and retrieve it thousands of tokens later.
Empirically, FwPKM demonstrates remarkable robustness. Models trained on short 4K-token sequences generalize effectively to 128K-token contexts (i.e. a $32\times$ length extrapolation) without fine-tuning. Furthermore, we show that iterative reading exploits the TTT nature of FwPKM, boosting retrieval accuracy from $10\%$ to $70\%$ in difficult Needle-in-a-Haystack scenarios. Our ablation study and interpretability analyses confirm that FwPKM functions as a true episodic memory, selectively activating for novel entities and complementing the semantic knowledge stored in standard slow weights.
Despite these successes, challenges remain. The introduction of online updates introduces additional computational overhead, currently limited by the lack of optimized kernels for sparse updates. This motivates future systems work on hardware-aware sparse kernels to close the speed gap with dense baselines. On the modeling side, our continual learning experiments reveal that while FwPKM adapts rapidly, it struggles with long-term retention across domain shifts. Future research should focus on advanced retention mechanisms and memory consolidation strategies to prevent catastrophic forgetting of fast weights. Overall, FwPKM represents a promising step toward language models with versatile and mutually complementary memory systems.
@@ Acknowledgments
We would like to thank Kai Arulkumaran, Luke Darlow, Stefania Druga, and the entire Sakana AI team for their insightful discussions and support throughout this project. We also extend our gratitude to Huayang Li, Qianying Liu, Richard Sproat, Yujin Tang, Qi Sun, and Stefano Peluchetti for their valuable feedback, which significantly strengthened the quality and clarity of this manuscript.
Appendix A FwPKM Implementation Details
A.1 Lookahead Value Targets
By default, FwPKM pairs each query $\mathbf{q}_{t}$ with the value computed from the same token, $\mathbf{v}_{t}$, in the memorization loss (Eq. 13). However, next-token prediction benefits more directly from information about future tokens. We therefore use a lookahead construction that pairs $\mathbf{q}_{t}$ with the next token’s value $\mathbf{v}_{t+1}$ when applying chunk-level updates.
Concretely, we keep the retrieval definition $\hat{\mathbf{v}}_{t}=\mathtt{PKM}(\mathbf{q}_{t};\theta)$ unchanged, and only shift the target value in the memorization objective. We use gradients from this new objective to update fast-weight value matrix $V$.
$\mathcal{L}_{\text{mem}}^{\text{LA}}$ $=\sum_{t=1}^{C-1}\frac{1}{2}\,g_{t}\,\|\mathbf{v}_{t+1}-\hat{\mathbf{v}}_{t}\|_{2}^{2}.$ (19)
Intuitively, this encourages FwPKM to store information that is immediately useful for predicting the next token, similar in spirit to using short convolutions in linear-attention variants [RWKV, GLA, Mamba].
Importantly, we apply the fast-weight update only after processing the entire chunk, so lookahead targets do not affect predictions within the same chunk. We carefully handle values at chunk boundaries. For efficiency, we drop the last predicted value in each chunk for efficiency at training time (which results in dropping 8 tokens for a 4096-token training sequence). At inference time, a FwPKM maintains a value cache that accumulates new $(\mathbf{v}_{t+1},\hat{\mathbf{v}}_{t})$ pairs. When a pre-defined update chunk size is reached (e.g. 512 in PPL evaluation, haystack length in NIAH evaluation), FwPKM consumes stored pairs to update its fast weights.
In our experiments, lookahead targets improved performance and we use it as the default unless otherwise noted.
A.2 Inverse-Distance Weighting (IDW) Scoring
A query–key score in PKM is typically the dot product $s_{i}=\mathbf{q}^{\top}K_{i}$. However, a key row can increase its score by growing its magnitude, without necessarily being close to the query in representation space. We therefore consider inverse distance weighting (IDW, IDW) as a drop-in alternative scoring function:
$s^{\text{IDW}}_{i}=-\log\big(\epsilon+\|\mathbf{q}-K_{i}\|_{2}^{2}\big),$ (20)
where we use $\epsilon=10^{-3}$. Due to the use of Euclidean distance, gradients produced by IDW scores push keys to behave as prototypes—centroids of query clusters. We apply IDW by replacing only the query–key scoring function inside PKM’s Top-$k$ selection (for each sub-key set in product-key retrieval); all other steps of PKM retrieval remain unchanged. Empirically, we found IDW scoring to yield better performance than dot-product scoring in our FwPKM setup.
A.3 Target Normalization and Gradient Clipping
We z-score normalize target values $\mathbf{v}_{t}$ along the feature dimension to stabilize optimization. Unlike standard slow-weight training, we do not clip gradients for fast-weight updates; this helps FwPKM match the scale of unbounded target values.
A.4 Loss Reduction and Effective Step Size
In Eq. 13 we sum over tokens and features. We avoid mean reduction so that the update magnitude does not shrink with chunk size $C$ or value dimension $d_{v}$. Concretely, if one uses mean reduction over $(t,j)$, the gradient of an element $V_{i,j}$ scales by $1/(Cd_{v})$, which effectively changes the fast-weight step size.
A.5 Gating-weighted Memorization
We weight each per-token MSE by $g_{t}$ in Eq. 13. Intuitively, $g_{t}$ measures how much the model relies on the memory output at this position, so it also serves as a natural importance weight for deciding which tokens should write more strongly. Note that we use gate $g_{t}$ both to interpolate outputs (Eq. 12) and to weight memorization updates in $\mathcal{L}_{\text{mem}}$ (and $\mathcal{L}_{\text{mem}}^{\text{LA}}$).
Appendix B Pseudo Code
1
2
Input : Hidden states $\mathbf{h}_{1:T}\in\mathbb{R}^{T\times d}$,
Fast weights $\theta=\{V\in\mathbb{R}^{N\times d_{v}},K^{(1)},K^{(2)}\in\mathbb{R}^{\sqrt{N}\times d_{k}/2}\}$,
Slow weights $\phi$
Hyperparameters : Chunk size $C$, Key dim $d_{k}$, Value dim $d_{v}$, Top-K $k$
Output : Output sequence $\mathbf{o}_{1:T}\in\mathbb{R}^{T\times d_{v}}$
3
4for $t=1$ to $T$ step $C$ do
 // --- Phase 1: Forward Pass (Inference) ---
 $\mathbf{h}_{\text{chunk}}\leftarrow\mathbf{h}_{t:t+C}$
 // Shape: $(C,d)$
5
 // Compute projections using Slow Weights $\phi$
 $\mathbf{q}\leftarrow\mathtt{Linear}^{q}_{\phi}(\mathtt{RMSNorm}^{q}_{\phi}(\mathbf{h}_{\text{chunk}}))$
 // Shape: $(C,d_{k})$
 $\mathbf{v}\leftarrow\mathtt{Linear}^{v}_{\phi}(\mathtt{RMSNorm}^{v}_{\phi}(\mathbf{h}_{\text{chunk}}))$
 // Shape: $(C,d_{v})$
 $g\leftarrow\sigma\big(\mathtt{Linear}^{g}_{\phi}(\mathtt{RMSNorm}^{g}_{\phi}(\mathbf{h}_{\text{chunk}}))\big)$
 // Shape: $(C,1)$
6
 // Split query into two halves for Product Keys
 $\mathbf{q}^{(1)},\mathbf{q}^{(2)}\leftarrow\mathtt{Split}(\mathbf{q})$
 // Shapes: $(C,d_{k}/2)$
7
 // Sparse Retrieval (PKM) with current Fast Weights
8 for $m\in\{1,2\}$ do
    // IDW Scoring
[…]
    // Shape: $(C,\sqrt{N})$
    // Top-k Indices
    $\mathcal{I}^{(m)}\leftarrow\mathtt{Top-k}(s^{(m)})$
    // Shape: $(C,k)$
9
10
 // Select final Top-k from the $k\times k$ restricted product
 $S_{\text{target}}\leftarrow\{s_{i}^{(1)}+s_{j}^{(2)}\mid(i,j)\in\mathcal{I}^{(1)}\times\mathcal{I}^{(2)}\}$
 // $k^{2}$ candidates per token
 $\mathcal{I}\leftarrow\mathtt{Top-k}(S_{\text{target}})$
 // Final $k$ indices per token
 $\hat{\mathbf{v}}\leftarrow\sum_{(i,j)\in\mathcal{I}}\mathtt{softmax}(S_{\text{target}})_{(i,j)}V_{\mathtt{flat-idx}(i,j)}$
 // Shape: $(C,d_{v})$
11
 // Gated Residual Output
 $\mathbf{o}_{\text{chunk}}\leftarrow g\cdot\hat{\mathbf{v}}+(1-g)\cdot\mathbf{v}$
 // Shape: $(C,d_{v})$
12
13
  1ex// --- Phase 2: Backward Pass for Memorization ---
 $\mathbf{v}_{\text{target}}\leftarrow\mathtt{Shift}(\mathbf{v},+1)$
 // Lookahead Target: $\mathbf{v}_{t+1}$
[…]
15
 $\nabla^{\text{agg}}_{V_{i}}\leftarrow\frac{1}{N_{i}^{\text{read}}}\nabla_{V_{i}}\mathcal{L}_{\text{mem}}$
 // Aggregate gradients by slot usage
 $V\leftarrow V-\nabla^{\text{agg}}_{V}$
 // Update Values
16
17
  1ex// --- Phase 3: Backward Pass for Addressing (Anti-collapsing) ---
18 for $m\in\{1,2\}$ do
    $\bar{p}^{(m)}\leftarrow\mathtt{AvgSlotUsage}(\mathcal{I}^{(m)},\text{chunk})$
    // Shape: $(\sqrt{N},)$
    $\mathcal{L}_{\text{addr}}^{(m)}\leftarrow-H(\bar{p}^{(m)})$
    // Maximize Entropy
    $K^{(m)}\leftarrow K^{(m)}-\nabla_{K^{(m)}}\mathcal{L}_{\text{addr}}^{(m)}$
    // Update Keys
19
20
return $\mathbf{o}_{1:T}$
Algorithm 1 Fast-weight Product Key Memory (FwPKM) Processing
Appendix C Detailed Experiment Settings